OpenStudy (anonymous):

What do you think about this code for problem 2?

OpenStudy (anonymous):

nugget_pacs=(input("Package Size 1 "),input("Package Size 2 "),input("Package Size 3 ")) # tuople contains nugget package sizes tracker=[] # list which will hold possible combinations not_tracker=[] # list to hold combinations not possible for n in range(1,150): # range of numbers to consider as candidates can_buy=0 for a in range(0,(n/nugget_pacs[0])+1): for b in range(0,(n/nugget_pacs[1])+1): for c in range(0,(n/nugget_pacs[2])+1): if nugget_pacs[0]*a+nugget_pacs[1]*b+nugget_pacs[2]*c ==n: can_buy=1 tracker.append(n) # adds the possible combo to list if tracker.count(n)>1: # if statement to remove repeats tracker.remove(n) if can_buy==0: not_tracker.append(n) # adds value of n not possible to list print 'Possible Combinations are' ,tracker print 'These Combinations are not possible' , not_tracker

OpenStudy (anonymous):

does it run? I do not get where you are taking nugget_pacs from?

OpenStudy (anonymous):

woops mom

OpenStudy (anonymous):

looks alright...if it runs an spits out good results it should be fine

OpenStudy (anonymous):

# chickennuggets.py # dio equations ### ### template of code for Problem 4 of Problem Set 2, Fall 2008 ### bestSoFar = 0 # variable that keeps track of largest number # of McNuggets that cannot be bought in exact quantity pack = (6,9,20) # variable that contains package sizes lim=55 lima=lim/6 limb=lim/9 limc=lim/20 solution=False list=[] for n in range(1, lim+1): # only search for solutions up to size 150 ## complete code here to find largest size that cannot be bought ## when done, your answer should be bound to bestSoFar solution=False for a in range(lima+1): for b in range(limb+1): for c in range(limc+1): trialn=pack[0]*a+pack[1]*b+pack[2]*c if trialn==n: solution=True break if solution: break if solution: break if solution: savenws=n list.append(n) print n, "has a solution, being", a, b, c if len(list)==6: break else: print n, "does not have a solution" saven=n list=[] print saven, "largest without solution, given package size", pack But I already incorporated that sequence of six from the next assignment...but apart from that it is similar in logic...

OpenStudy (anonymous):

also I did not save all possible combos...I took the first best and left

OpenStudy (anonymous):

nugget_pacs is a tuople containing the size of the different packages which is taken in as input from the user. It then callc the specific package size by calling its index